A single force of 40 N acts upon a 5 kg block. what is the magnitude of the acceleration of the block?
1
Expert's answer
2012-09-18T12:13:33-0400
Let's use the second Newton's law:

F = ma ==> a = F/m = 40/5 = 8 m/s².

So, the magnitude of the acceleration of the block is 8 m/s².